OpenStudy (anonymous):

Does the data in the table represent a direct variation or an inverse variation? Write an equation to model the data in the table. x 2 4 8 12 y 4 2 mc019-1.jpg 2/3 direct variation; y = 8/x direct variation; y = 8x inverse variation; xy = 8 inverse variation; y/x = 8

OpenStudy (anonymous):

@LynFran

OpenStudy (anonymous):

@AngusV

OpenStudy (anonymous):

@LynFran

OpenStudy (lynfran):

whats mc019.jpg

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

I'm stumped on this also

OpenStudy (lynfran):

y=kx is direct variation y=k/x is an inverse variation

OpenStudy (lynfran):

so lets see

OpenStudy (lynfran):

when x=2 y=4 when x=4 y=2 when x=8 y=?? when x=12 y=2/3

OpenStudy (lynfran):

whats y when x=8

OpenStudy (lynfran):

im not seeing the value

OpenStudy (lynfran):

this is what im seeing y 4 2 mc019-1.jpg =====whats this value??? 2/3

OpenStudy (anonymous):

it is 1

OpenStudy (lynfran):

ok

OpenStudy (lynfran):

y=kx if its direct variation we will use when x=2 y=4 so 4=k2 4/2=k 2=k we will use when x=4 y=2 so 2=k4 2/4=k 1/2=k when x=8 y=1 1=k8 1/8=k as we see k is constantly changing so...this is not direct variation

OpenStudy (anonymous):

so what is it?

OpenStudy (lynfran):

y=k/x inverse variation when x=2 y=4 4=k/2 4*2=k 8=k when x=4 y=2 2=k/4 2*4=k 8=k when x=8 y=1 1=k/8 1*8=k 8=k so yes its inverse variation k=8

OpenStudy (lynfran):